OpenStudy (anonymous):

I need help on converting Slope Intercept form into Standard form. y=2x-0

OpenStudy (anonymous):

does it really say \(y=2x-0\) ???

OpenStudy (anonymous):

Yes, it needs to be converted to Standard form

OpenStudy (anonymous):

most humans would just write \(y=2x\)

OpenStudy (anonymous):

subtract \(y\) and get \[2x-y=0\] in standard form

OpenStudy (anonymous):

yes I know, but how do I get there

OpenStudy (pinksapphire):

y-y1=m(x-x1)

OpenStudy (anonymous):

just subtract y?

OpenStudy (pinksapphire):

Pick any coordinate that will agree to the equation...

OpenStudy (pinksapphire):

Like (2,4)

OpenStudy (anonymous):

Well I put the coordinates into Slope-Intercept into Point-Slope..

OpenStudy (anonymous):

was that a mistake?

OpenStudy (anonymous):

Instead of Point-Slope*

OpenStudy (pinksapphire):

y-4=2x(x-2)

OpenStudy (anonymous):

yes, thank you. I think I found my answer.

OpenStudy (pinksapphire):

Did you want the standard form?

OpenStudy (anonymous):

well I wanted the steps to the standard form

OpenStudy (anonymous):

not just the answer

OpenStudy (anonymous):

I will give you a medal if you give me the steps

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@Mindblast3r

OpenStudy (pinksapphire):

Ax+By=C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y-4=2(x-2) y-4=2x-4

OpenStudy (anonymous):

then what after?

OpenStudy (pinksapphire):

No, that is the point-slope form.

OpenStudy (anonymous):

yes I understand. How do I now get Point-Slope into Standard form?

OpenStudy (anonymous):

I am now rewriting the whole thing

OpenStudy (pinksapphire):

y=2x+0 is your slope-intercept form and we are going to make it into standard form (Ax+By=C)

OpenStudy (anonymous):

with all of the steps?

OpenStudy (anonymous):

if so, that would be much appreciated.

OpenStudy (pinksapphire):

Do you want a video to explain it?

OpenStudy (anonymous):

I just need steps. I have looked at a lot of videos

OpenStudy (anonymous):

if you make the video then sure

OpenStudy (pinksapphire):

Maybe this video will help: https://www.youtube.com/watch?v=JeR_fuO__2M

OpenStudy (anonymous):

you cannot do a quick run through of the steps?

OpenStudy (pinksapphire):

I'll try...

OpenStudy (anonymous):

thank you

OpenStudy (pinksapphire):

y=2x+0 Start with your slope-intercept form. y=2x+0 Subtract 2x on both sides. -2x -2x+y=0 This is your standard form equation.

OpenStudy (anonymous):

thanks, i'll give your medal

OpenStudy (pinksapphire):

Thanks :)